Grade 12Integral Calculus

evaluate
∫ sqrt(x){3 – 5x} dx

taniska
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \sqrt{x}(3-5x)dx
I = \int (3\sqrt{x}-5x^{3/2})dx
I = \int 3\sqrt{x}dx-\int 5x^{3/2}dx
I = 3.\frac{x^{3/2}}{3/2}- 5.\frac{x^{5/2}}{5/2}
I = 2x^{3/2} - 2x^{5/2}+c